# SADDE: Semi-supervised Anomaly Detection with Dependable Explanations

### v1

@@
Abstract.
Semi-supervised learning holds a pivotal position in anomaly detection applications, yet identifying anomaly patterns with a limited number of labeled samples poses a significant challenge. Furthermore, the absence of interpretability poses major obstacles to the practical adoption of semi-supervised frameworks. The majority of existing interpretation techniques are tailored for supervised/unsupervised frameworks or non-security domains, falling short in providing dependable interpretations. In this research paper, we introduce SADDE, a general framework designed to accomplish two primary objectives: (1) to render the anomaly detection process interpretable and enhance the credibility of interpretation outcomes, and (2) to assign high-confidence pseudo labels to unlabeled samples, thereby boosting the performance of anomaly detection systems when supervised data is scarce. To achieve the first objective, we devise a cutting-edge interpretation method that utilizes both global and local interpreters to furnish trustworthy explanations. For the second objective, we conceptualize a novel two-stage semi-supervised learning framework tailored for network anomaly detection, ensuring that the model predictions of both stages align with specific constraints. We apply SADDE to two illustrative network anomaly detection tasks and conduct extensive evaluations in comparison with notable prior works. The experimental findings underscore that SADDE is capable of delivering precise detection results alongside dependable interpretations for semi-supervised network anomaly detection systems. The source code for SADDE is accessible at: https://github.com/M-Code-Space/SADDE.
@@ Abstract.
Dependable explanation, semi-supervised learning, anomaly detection
1. Introduction
@@
Anomaly detection has become a fundamental task in various applications, including network intrusion detection (Mirsky et al., 2018), web attack detection (Tang et al., 2020), and advanced persistent threat detection (Bowman et al., 2020). Anomaly detection systems aim to identify unforeseen threats, such as zero-day attacks. To achieve this, semi-supervised learning is increasingly promising, as it requires only limited labeled data, unlike traditional supervised approaches that depend on extensive labeled samples for training.
@@
Despite the great promise of semi-supervised network anomaly detection systems, the lack of interpretability of their predictions poses major barriers to their practical adoption. Firstly, it is challenging to build trust in the decisions made by these systems, which often provide only binary outputs (benign or malicious) without adequate justification or reliable evidence. Second, it is nearly impossible for analysts to manually interpret the decisions of these black-box machine learning models, as the volume of training data and the complexity of the learned models are beyond human understanding (Guidotti et al., 2018b). Third, minimizing false positives and miss-detected samples remains a critical challenge for anomaly detection systems in real-world applications. Without a clear understanding of how the models operate, it is impossible to effectively update or adjust them to reduce them. Consequently, security operators are left unable to decide whether they can trust model decisions based on excessively simple model predictions, are unwilling to use theoretical semi-supervised network anomaly detection systems in their practical applications, and feel overwhelmed by numerous meaningless false positives.
Figure 1. Illustration of the divergent explanation results of different interpreters.
Figure 2. Overview of the proposed SADDE.
@@
Recently, some research has proposed techniques for interpreting machine learning model decisions by identifying key features that significantly influence the final prediction (Montavon et al., 2018; Ribeiro et al., 2016b; Lundberg and Lee, 2017; Guo et al., 2018). Predominant approaches include local/model-agnostic (Ribeiro et al., 2016a, 2018; Guidotti et al., 2018a; Lundberg, 2017; Guo et al., 2018) and global (Fisher et al., 2019; Kauffmann et al., 2024) methods that focus on interpreting individual and clusters predictions for a given black-box classifier. These interpreters either try to approximate the local decision boundary using a linear algorithm to find key features of the given input, or they perturb the current input’s feature values or permute features of a cluster and observe the model’s feedback to pinpoint the most influential ones. Security operators can use these interpreted key features to better understand model decisions (Guidotti et al., 2018b). Existing work like (Fong and Vedaldi, 2017; Shrikumar et al., 2017a; Zhou et al., 2016; Zintgraf et al., 2017; Kauffmann et al., 2024; Mothilal et al., 2020; Shrikumar et al., 2017b) predominantly focuses on interpreting supervised/unsupervised models in non-security domains, while some existing interpreters like (Han et al., 2021; Liu et al., 2017a; Yang et al., 2021; Wei et al., 2023) are proposed for security-related applications. However, these studies cannot be directly used for semi-supervised network anomaly detection. Furthermore, studies indicate that the explanation results provided by existing interpreters lack consensus (Fan et al., 2020), leading to confusion regarding their trustworthiness. To our best knowledge, although a few studies examine the reliability of existing interpreters (Fan et al., 2020; Yeh et al., 2019; Camburu et al., 2019), they mainly focus on developing more effective evaluation metrics rather than improving the reliability of interpreters’ results. Fig. 1 illustrates the process of a classifier analyzing network traffic data and producing decisions, such as identifying whether the traffic is anomalous or benign. Various interpretability methods, including LIME, DeepLIFT, Lemna, and so on, generate different key features (e.g., ”Flow Bytes/s,” ”Idle Max,” ”Fwd IAT Min”), leading to confusion for the security analyst. The inconsistency in these explanations raises uncertainty regarding the reliability of the interpretations.
@@
Our Work. In this paper, we propose SADDE, a general semi-supervised framework for automatically learning from enormous unlabeled data and improving the reliability of interpretation results. The design goal of SADDE is to develop a novel semi-supervised learning framework for network anomaly detection applications that meets the special requirements of security domains (such as reliable, human-understandable, stable, robust, and fast). To this end, we propose two techniques in SADDE referred to as Global-local Knowledge Association Mechanism (KAM) and Two-stage Semi-supervised Learning System (ToS). KAM helps security analysts understand the underlying reason for model predictions while ToS provides an effective pseudo-labeling process to assist the semi-supervised learning. Compared to original black-box machine learning models, security practitioners can better understand system feedback. Besides, with more reliable interpretations, system operators are more willing to adopt theoretical semi-supervised frameworks in their practical applications. We also provide prototype implementations of SADDE over two representative network anomaly detection datasets (ISCXTor2016 (Lashkari et al., 2017) and CIC-DoHBrw-2020 (MontazeriShatoori et al., 2020)). The proposed KAM and ToS are extensively evaluated compared to representative prior approaches. Experimental results demonstrate that SADDE can provide reliable explanations for semi-supervised learning frameworks while satisfying the requirements of security-related applications.
@@
  • We propose SADDE, a general framework for interpreting and improving semi-supervised network anomaly detection in security-related applications, which includes two key components: KAM provides reliable, human-understandable, stable, robust, and fast interpretations for semi-supervised systems and ToS makes it possible to automatically learn from large unlabeled datasets.
@@
  • We conduct extensive experiments to demonstrate that the proposed SADDE outperforms existing approaches regarding fidelity, stability, robustness, and efficiency.
2. Framework
2.1. Overview
@@
The overview of SADDE is illustrated in Fig. 2 consisting of two key components: Global-local Knowledge Association Mechanism (KAM) and Two-stage Semi-supervised Learning System (ToS). ToS enables accurate semi-supervised learning via a novel two-stage pseudo-labeling process while KAM provides reliable interpretations for the semi-supervised anomaly detection system. SADDE is a general model-independent framework that formulates semi-supervised learning and decision interpretation into a unified end-to-end learning process with security-related constraints. We instantiate SADDE on the aforementioned network anomaly detection security application. With SADDE, malicious network traffic can be automatically detected without requiring large labeled datasets, meanwhile, security practitioners can better understand system behaviors and trust system decisions.
@@
Working Process. The working process of the proposed SADDE is as follows. Before starting, we assume that there is a small labeled dataset that can be utilized to pre-train the local and global machine learning models. Such assumption is reasonable, as semi-supervised learning necessitates the model to continuously assign pseudo-labels to unlabeled samples, and pre-training the model using a small labeled dataset saves data labeling costs while ensuring the effectiveness of the subsequent ”data-labeling” process. Similar assumptions can be found in (Yao et al., 2018; Apruzzese et al., 2022; Li et al., 2024). Given an instance (or a batch of instances), we fed it into the pre-trained local model to extract its feature embedding and predict its label. The extracted feature embedding is then clustered by the global model into several clusters, where samples from the same cluster share similar characteristics. Feature representations from a cluster are treated as global representations while the feature embedding from an instance is defined as a local representation. We group the input instance into one of the clusters by comparing the distance between the embedding of the instance and the clusters. Then, the predictions of the global and local models are interpreted by the global and local interpreters, and the reliability of the interpretation results is further improved by comparing the global and local interpretations. The underlying reason behind it is that global and local interpreters are heterogeneous interpreters. It has a higher probability of improving its trustworthiness by matching global and local interpretations instead of two local interpreters. For instance, in scenarios where individual samples contain noise or have been maliciously altered, the interpretation results at the local level may be wrong, however, due to the fact that the global and local interpreters are heterogeneous, there is a probability that the interpretation results at the global level is correct. In this way, the interpretations’ reliability could be improved. If the interpreted global and local key features match (the similarity score is higher than a given threshold), we consider the key features interpreted by the interpreters to be credible. Conversely, if they do not match, the samples will be discarded. These samples will not be utilized for model re-training to avoid possible model performance deterioration. Finally, labels predicted by the global and local models are aligned to find the most credible pseudo-labels for the next round’s semi-supervised training. More details about SADDE are presented in Algorithm 1.
@@
             Calculate a similarity score $sim$ between the interpreted global and local key features by
@@
Algorithm 1 SADDE
2.2. Global-local Knowledge Association Mechanism (KAM)

### v4

# AnomalyAID: Reliable Interpretation for Semi-supervised Network Anomaly Detection

@@
Abstract
Semi-supervised Learning plays a crucial role in network anomaly detection applications, however, learning anomaly patterns with limited labeled samples is not easy. Additionally, the lack of interpretability creates key barriers to the adoption of semi-supervised frameworks in practice. Most existing interpretation methods are developed for supervised/unsupervised frameworks or non-security domains and fail to provide reliable interpretations. In this paper, we propose AnomalyAID, a general framework aiming to (1) make the anomaly detection process interpretable and improve the reliability of interpretation results, and (2) assign high-confidence pseudo labels to unlabeled samples for improving the performance of anomaly detection systems with limited supervised data. For (1), we propose a novel interpretation approach that leverages global and local interpreters to provide reliable explanations, while for (2), we design a new two-stage semi-supervised learning framework for network anomaly detection by aligning both stages’ model predictions with special constraints. We apply AnomalyAID over two representative network anomaly detection tasks and extensively evaluate AnomalyAID with representative prior works. Experimental results demonstrate that AnomalyAID can provide accurate detection results with reliable interpretations for semi-supervised network anomaly detection systems.
keywords:
@@ Abstract
Explainable machine learning , Reliable explanations , Semi-supervised learning , Network anomaly detection
22footnotetext: Corresponding author(s): Yingwen Wu and Jin Wang
[…]
[1]organization=School of Future Science and Engineering, Soochow University,city=Suzhou, state=Jiangsu, country=China
[…]
[3]organization=Southeast University, city=Nanjing, state=Jiangsu, country=China
1 Introduction
@@
Anomaly detection has become a fundamental task in various applications, including network intrusion detection [mirsky2018kitsune], web attack detection [9155278], and advanced persistent threat detection [259729]. Anomaly detection systems aim to identify unforeseen threats, such as zero-day attacks. To achieve this, semi-supervised learning is increasingly promising, as it requires only limited labeled data, unlike traditional supervised approaches that depend on extensive labeled samples for training.
@@
Despite the great promise of semi-supervised network anomaly detection systems, the lack of interpretability of their predictions poses major barriers to their practical adoption. Firstly, it is challenging to build trust in the decisions made by these systems, which often provide only binary outputs (benign or malicious) without adequate justification or reliable evidence. Second, it is nearly impossible for analysts to manually interpret the decisions of these black-box machine learning models, as the volume of training data and the complexity of the learned models are beyond human understanding [guidotti2018survey]. Third, minimizing false positives and miss-detected samples remains a critical challenge for anomaly detection systems in real-world applications. Without a clear understanding of how the models operate, it is impossible to effectively update or adjust them to reduce them [han2021deepaid]. Consequently, security operators are left unable to decide whether they can trust model decisions based on excessively simple model predictions, are unwilling to use theoretical semi-supervised network anomaly detection systems in their practical applications, and feel overwhelmed by numerous meaningless false positives.
Figure 1: Illustration of the divergent explanation results of different interpreters.
@@
Recently, some research has proposed techniques for interpreting machine learning model decisions by identifying key features that significantly influence the final prediction [ribeiro2016should_lime, lundberg2017unified_shap, guo2018_lemna]. Predominant approaches include local/model-agnostic [ribeiro2016should, ribeiro2018anchors, lundberg2017unified_shap, guo2018_lemna] and global [fisher2019all, 9817459] methods that focus on interpreting individual and cluster predictions for a given black-box classifier. These interpreters either try to approximate the local decision boundary using a linear algorithm to find key features of the given input, or they perturb the current input’s feature values or permute features of a cluster and observe the model’s feedback to pinpoint the most influential ones. Security operators can use these interpreted key features to better understand model decisions [guidotti2018survey]. Existing work like [fong2017interpretable, shrikumar2017learning, 9817459, mothilal2020explaining, shrikumar2017learning_deeplift, han2021deepaid, wei2023xnids] predominantly focuses on interpreting supervised/unsupervised models in various applications. However, the field of developing reliable interpretation for semi-supervised network anomaly detection remains unexplored. Fig. 1 illustrates the process of a classifier analyzing network traffic data and producing decisions, such as identifying whether the traffic is anomalous or benign. Various interpretability methods, including LIME [ribeiro2016should_lime], DeepLIFT [shrikumar2017learning_deeplift], and Lemna [guo2018_lemna], generate different key features (e.g., ”Flow Bytes/s,” ”Idle Max,” and ”Fwd IAT Min”), leading to confusion for the security analyst. The inconsistency in these explanations raises uncertainty regarding the reliability of the interpretations.
@@
Our Work. In this paper, we propose AnomalyAID, a general semi-supervised framework for automatically learning from enormous unlabeled data and improving the reliability of interpretation results. The design goal of AnomalyAID is to develop a novel semi-supervised learning framework for network anomaly detection applications that meets the special requirements of security domains (such as reliability). To this end, we propose two techniques in AnomalyAID referred to as Global-local Knowledge Association Mechanism (KAM) and Two-stage Semi-supervised Learning System (ToS). KAM helps security analysts understand the underlying reason for model predictions, while ToS provides an effective pseudo-labeling process to assist the semi-supervised learning. Compared to original black-box machine learning models, security practitioners can better understand system feedback. Besides, with more reliable interpretations, system operators are more willing to adopt theoretical semi-supervised frameworks in their practical applications. We also provide prototype implementations of AnomalyAID over two representative network anomaly detection datasets (ISCXTor2016 [lashkari2017characterization], CIC-DoHBrw-2020 [montazerishatoori2020detection], and UNSW-NB15 [moustafa2015unsw]). The proposed KAM and ToS are extensively evaluated with representative prior approaches. Experimental results demonstrate that AnomalyAID can provide reliable explanations for semi-supervised learning frameworks while satisfying the requirements of security-related applications (Our code is available at: https://github.com/M-Code-Space/AnomalyAID.).
@@
    We propose AnomalyAID, a general framework for interpreting and improving semi-supervised network anomaly detection in security-related applications, which includes two key components: KAM provides reliable interpretations for semi-supervised systems and ToS makes it possible to learn from large unlabeled datasets automatically.
  • 2.
@@
    We conduct extensive experiments to demonstrate that the proposed AnomalyAID outperforms existing approaches regarding fidelity, stability, robustness, and efficiency.
2 Motivating Scenarios
For network anomaly detection applications, a security analyst relies on a machine learning-based algorithm to identify network anomalies. The machine learning model is trained in a supervised or semi-/un-supervised manner. After training the detector, the security analyst analyzes incoming samples and classifies them as either malicious or benign. However, since the model operates as a black box, it doesn’t provide any reasoning behind its decisions. To understand why the system flagged a sample, the analyst turns to explainability methods that highlight critical features of the input. Fig. 2 shows the interpretation outcomes for an input sample identified as anomalies by the same detector, employing six distinct local interpreters: LIME, DeepLift, SHAPE, COIN, CADE, and xNIDS.
Figure 2: Six example explanation results for the same input data with the same model using LIME, DeepLift, SHAPE [lundberg2017unified_shap], COIN [liu2017contextual_coin], CADE [yang2021cade], and xNIDS [wei2023xnids]. Different interpreters give different explanations.
In Fig. 2, the interpretation results of LIME [ribeiro2016should_lime], DeepLift [shrikumar2017learning_deeplift], SHAPE [lundberg2017unified_shap], COIN [liu2017contextual_coin], CADE [yang2021cade], and xNIDS [wei2023xnids] highlight five key features (when the security analyst sets the number of key features as five). Note that the numerical values associated with these features represent their respective impact weights. Interestingly, the outputs from the six methods exhibit significant differences. For instance, LIME and DeepLift show no overlapping features in their results. Additionally, while the Source IP feature in SHAP has a negative influence on the classification of an input network traffic instance, in COIN, it demonstrates an opposing effect. A similar phenomenon can be observed for the Source Port feature. These discrepancies leave the analyst perplexed. He/she may wonder which method to select and whether these explanations can be relied upon. Given the lack of standardized metrics for assessing explainability techniques in network anomaly detection, this situation underscores the need for an evaluation study to assess the reliability and applicability of these methods in this critical area.
Network anomaly detection. In the process of detecting network anomalies (i.e., classifying network traffic into malicious or benign traffic), each network traffic flow $x_{i}$ is represented as a feature vector in a $d$-dimensional space, where $x_{i}\in\mathbb{R}^{d}$. A network anomaly detector $f$ can be developed using a designated supervised or semi-/un-supervised machine learning algorithm based on this dataset. We can express the prediction for the $i$-th sample as $f(x_{i})=\hat{y}_{i}$, where $\hat{y}_{i}$ reflects the predicted output associated with $x_{i}$. The classifier is considered to perform correctly if the predicted label $\hat{y}_{i}$ aligns with the actual label $y_{i}$.
Interpretation. The interpretation process is denoted as $g=m(f)$, where $f$ is a specific pre-trained network anomaly detector and $m$ is an interpreter. Given a network traffic flow $x_{i}$, the interpreter $m$ will produce the interpretation result $g$, which is represented as a list of features that are ranked based on their contribution to the final decision.
3 Preliminaries
3.1 Explanation methods
PFIE. PFIE, proposed by [fisher2019all], is a simple and easy-to-implement feature importance calculation method used to explain the predictions of machine learning models. Its core idea is to evaluate the importance of a feature to the model’s predictions by observing the change in the model’s prediction error after shuffling (or permuting) the values of that feature. Specifically, the calculation of PFIE first involves computing the baseline error using the trained model and the original dataset. Then, a specific feature is selected, and its values are randomly shuffled to break the relationship between the feature and the target variable. Next, the model’s prediction error is recalculated using the shuffled dataset. Finally, the importance of the feature is assessed by comparing the baseline error with the shuffled error (as defined in Eq., (1)). The greater the increase in error, the more important the feature is to the model’s predictions.
$\textit{Importance}(f)=\textit{Error}(M)-\textit{Error}(M_{p}),$ (1)
where $\textit{Importance}(f)$ means the permutation importance of feature $f$, indicating the impact of the feature $f$ on the model’s performance. The larger the value, the more important the feature. $\textit{Error}(M)$ is the error metric of model $M$ on the original data, representing the model’s performance without modifying the feature. $\textit{Error}(M_{p})$ denotes the error metric of the shuffled model $M_{p}$ on the dataset, reflecting the change in model performance after shuffling feature $f$.
DeepLift. It is introduced by [shrikumar2017learning_deeplift], which is designed to interpret the output predictions of neural networks by comparing the activation of each neuron to its “reference activation” and decomposing the contributions of all neurons in the network to the input features through backpropagation. In the context of network traffic anomaly detection, the process begins with selecting a zero-input as the reference input to compute the difference in activations. Typically, the reference input is chosen to be a zero vector (i.e., all feature values are set to zero), representing a state of “no features.” For instance, $x_{0}=0$. Next, both the reference input $x_{0}$ and the actual input $x_{i}$ are propagated through the deep learning-based network traffic anomaly detection model to compute the reference activation $y_{0}$ and the actual activation $y_{i}$ for each neuron. The difference between the reference activations and the actual activations is then calculated for each neuron. Subsequently, the rules of DeepLift—such as the Linear Rule, Rescale Rule, or RevealCancel Rule—are applied to assign contribution scores for each input feature (e.g., pixel) to the output. Finally, the contribution scores are backpropagated from the output layer to the input layer using the chain rule. Through this approach, DeepLift provides a clear interpretation of how the CNN model utilizes input features for decision-making, thereby enhancing the model’s interpretability.
3.2 Machine learning-based classifiers
DBSCAN. Density-Based Spatial Clustering of Applications with Noise (DBSCAN) [ester1996density] is a popular clustering algorithm that focuses on identifying clusters based on the density of data points. Unlike traditional clustering methods like K-Means, DBSCAN does not require the number of clusters to be specified in advance and can discover clusters of arbitrary shapes while effectively handling noise.
Specifically, it calculates the density around each sample using two parameters: the neighborhood radius $\epsilon$ and the minimum number of points $MinPts$ required to form a dense region. The Euclidean distance between two samples $G(x_{i})$ and $G(x_{k})$ is denoted as $\|G(x_{i})-G(x_{k})\|$. The $\epsilon$-neighborhood of a sample $x_{i}$ is defined as:
$N_{\epsilon}(x_{i})=\{x_{j}\mid\|G(x_{i})-G(x_{j})\|\leq\epsilon\},$ (2)
where $N_{\epsilon}(x_{i})$ represents the set of points within the $\epsilon$-neighborhood of $x_{i}$. If the number of points in the neighborhood, $|N_{\epsilon}(x_{i})|$, is at least $MinPts$, the sample $x_{i}$ is classified as a core point. Samples reachable from a core point but not meeting the density requirement are labeled as border points, while samples that are not reachable from any core point are treated as outliers.
FCN. A Fully Connected Network (FCN) [rumelhart1986learning] is a type of neural network architecture where each neuron in one layer is connected to every neuron in the next layer. This architecture is commonly used for tasks such as network traffic anomaly detection. The key ideas and principles of an FCN revolve around the transformation of input data through multiple layers of neurons, with each layer learning increasingly abstract representations of the input.
An FCN consists of multiple layers, including input layers, hidden layers, and an output layer. Each layer is fully connected to the next layer, meaning every neuron in one layer is connected to every neuron in the subsequent layer (the calculation is shown in Eq. (3)).
$\mathbf{z}^{(l+1)}=\mathbf{W}^{(l)}\mathbf{x}^{(l)}+\mathbf{b}^{(l)},$ (3)
where $\mathbf{W}^{(l)}$ and $\mathbf{b}^{(l)}$ are the weight matrix and bias vector of layer $l$. $\mathbf{x}^{(l)}$ is the input to layer $l$. Each neuron in an FCN computes a weighted sum of its inputs and applies a non-linear activation function (e.g., ReLU) to introduce non-linearity into the model. This allows the network to learn complex patterns in the data. FCNs are trained using backpropagation, where the error between the predicted output and the true label is propagated backward through the network to update the weights. This process minimizes a loss function, usually cross-entropy for classification (see Eq. (4)).
$\mathcal{L}=-\sum_{i}t_{i}\log(y_{i}),$ (4)
where $t_{i}$ is the true label and $y_{i}$ is the predicted probability for class $i$.
In our study, we use the PFIE and DeepLift as our global and local interpreters, and FCN and DBSCAN as our local and global models, respectively, as an example for evaluating the effectiveness of the proposed ToS. Other machine learning models can also be chosen accordingly for specific tasks.
Figure 3: Overview of the proposed AnomalyAID.
4 Framework
4.1 Overview
@@
The overview of AnomalyAID is illustrated in Fig. 3, consisting of two key components: Global-local Knowledge Association Mechanism (KAM) and Two-stage Semi-supervised Learning System (ToS). ToS enables accurate semi-supervised learning via a novel two-stage pseudo-labeling process, while KAM provides reliable interpretations for the semi-supervised anomaly detection system. AnomalyAID is a general model-independent framework that formulates semi-supervised learning and decision interpretation into a unified end-to-end learning process with security-related constraints. We instantiate AnomalyAID on the aforementioned network anomaly detection security application. With AnomalyAID, malicious network traffic can be automatically detected without requiring large labeled datasets; meanwhile, security practitioners can better understand system behaviors and trust system decisions.
Working Process. The working process of the proposed AnomalyAID is as follows. Before starting, we assume that there is a small labeled dataset that can be utilized to pre-train the local and global machine learning models. Such an assumption is reasonable, as semi-supervised learning necessitates the model to continuously assign pseudo-labels to unlabeled samples, and pre-training the model using a small labeled dataset saves data labeling costs while ensuring the effectiveness of the subsequent data-labeling process. Similar assumptions can be found in [li2024instantT].
@@
Given an instance (or a batch of instances), we fed it into the pre-trained local model to extract its feature embedding and predict its label. The extracted feature embedding is then clustered by the global model into several clusters, where samples from the same cluster share similar characteristics. Feature representations (or feature embeddings, we use them interchangeably in this paper) from a cluster are treated as global representations, while the feature embedding from an instance is defined as a local representation. We group the input instance into one of the clusters by comparing the distance between the embedding of the instance and the clusters. Then, the predictions of the global and local models are interpreted by the global and local interpreters, and the reliability of the interpretation results is further improved by comparing the global and local interpretations. The underlying reason behind this is that global and local interpreters are heterogeneous interpreters. It has a higher probability of improving its trustworthiness by matching global and local interpretations instead of two local interpreters.
For instance, in scenarios where individual samples contain noise or have been maliciously altered, the interpretation results at the local level may be wrong; however, due to the fact that the global and local interpreters are heterogeneous, there is a probability that the interpretation results at the global level are correct. In this way, the interpretations’ reliability could be improved. If the interpreted global and local key features match (the similarity score is higher than a given threshold), we consider the key features interpreted by the interpreters to be credible. Conversely, if they do not match, the samples will be discarded. These samples will not be utilized for model re-training to avoid possible model performance deterioration. Finally, labels predicted by the global and local models are aligned to find the most credible pseudo-labels for the next round’s semi-supervised training. More details about AnomalyAID are presented in Algorithm 1.
@@
      Calculate a similarity score simsim between the interpreted global and local key features by
@@
Algorithm 1 AnomalyAID
4.2 Global-local Knowledge Association Mechanism (KAM)